# Univariate-Guided Sparse Regression

### v8

@@ 15 Discussion
In this paper we have introduced a novel method for sparse regression that “stacks” univariate regressions with a non-negative form of the Lasso. The procedure has interesting properties, namely that the final features weights have the same signs as the univariate coefficients and tend to be larger when the univariate coefficients are large. The test set MSE of the method tends to be similar to that of the Lasso, with substantially smaller support and lower false positive rate.
There are many possible extensions for this work. For example, rather than starting with raw features, one could apply UniLasso to derived features such as trees from random forests or gradient boosting. This is a topic for future study.
Draft versions of the UniLasso package can be found at
The Python package was written by Sophia Lu.
@@ Acknowledgments
The authors would like to thank Jerome Friedman, Alden Green, Sam Gross, Lucas Janson, Mert Pilanci, Jonathan Taylor, Ryan Tibshirani and Hui Zou for helpful discussions. We especially thank Ryan for posing and proving theorem 3., and Alden for the suggestion to investigate adversarial attacks in Section 10. S.C. was supported by NSF grants DMS-2113242 and DMS-2153654. T.H was supported by grant 5R01 EB 001988-21 from the National Institutes of Health; R.T. was supported by the NIH (5R01EB001988- 16) and the NSF (19DMS1208164).
Appendix
Proof of Theorem 1
We will need the following result (see, e.g., (?, Theorem A.7.1)).
@@
Choosing $\beta$ sufficiently small makes the upper bound $\leq 2$, allowing us to apply Berstein’s inequality. A similar argument works for $B_{j}$. For $D_{j}$, take $Z_{i}:=Y_{i}X_{i,j}-\rm E(D_{j})$. Then note that $\rm E(Z_{i})=0$, and for any $t\geq 2|\rm E(D_{j})|$,
@@
There are positive constants $C_{5}$, $C_{6}$ and $C_{7}$ depending only on $C_{0}$, $C_{1}$ and $C_{2}$ such that for any $t\in[0,C_{5}]$ and any $i$ and $j$, $\mathbb{P}(|\hat{\beta}_{i,j}-\beta_{j}|\geq t)$ and $\mathbb{P}(|\hat{\alpha}_{i,j}-\alpha_{j}|\geq t)$ are bounded above by $C_{6}e^{-C_{7}nt^{2}}$.
@@
$\mathbb{P}(|\hat{\beta}_{i,j}-\beta_{j}|\geq t)$ $=\mathbb{P}\biggl{(}\biggl{|}\frac{Q_{1}-Q_{2}Q_{3}-(Q_{4}-Q_{3}^{2})\beta_{j}}{Q_{4}-Q_{3}^{2}}\biggr{|}\geq t\biggr{)}$
@@
Choosing $s$ to solve $K_{5}s+K_{6}s^{2}=C_{0}t/2$, and combining equations (12), (14) and (15) yields the proof of the claimed concentration inequality for $\hat{\beta}_{i,j}$. The proof for $\hat{\alpha}_{i,j}$ is similar. We omit the details. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i,j}-\hat{\alpha}_{j})^{2}$ $\leq\frac{2}{n}\sum_{i=1}^{n}(\hat{Y}_{i,j}-\hat{\alpha}_{i,j})^{2}+\frac{2}{n}\sum_{i=1}^{n}(\hat{\alpha}_{i,j}-\hat{\alpha}_{j})^{2}$
$\leq\frac{2}{n}\sum_{i=1}^{n}\hat{\beta}^{2}_{i,j}X_{i,j}^{2}+\frac{4}{n}\sum_{i=1}^{n}(\hat{\alpha}_{i,j}-\alpha_{j})^{2}+4(\hat{\alpha}_{j}-\alpha_{j})^{2}$
$\leq\biggl{(}\frac{2}{n}\sum_{i=1}^{n}X_{i,j}^{2}\biggr{)}\max_{1\leq i\leq n}\hat{\beta}_{i,j}^{2}+4\max\{(\hat{\alpha}_{j}-\alpha_{j})^{2},\max_{1\leq i\leq n}(\hat{\alpha}_{i,j}-\alpha_{j})^{2}\}.$
@@
because otherwise, we can slightly perturb $\hat{\theta}_{j}$ while maintaining the nonnegativity constraint and decreasing the value of $L$. By inequality (17), this implies that if $\hat{\theta}_{j}>0$, then we must have
@@
$M:=\max_{1\leq i\leq n,\,j\in S}(|\hat{\beta}_{i,j}-\beta_{j}|+|\hat{\beta}_{j}-\beta_{j}|).$
@@
$|\hat{\theta}_{j}(\hat{Y}_{i,j}-\hat{\alpha}_{j})-\hat{\gamma}_{j}X_{i,j}|$ $=|\hat{\theta}_{j}\hat{\beta}_{i,j}X_{i,j}-\hat{\theta}_{j}\hat{\beta}_{j}X_{i,j}|$
$=|\hat{\beta}_{i,j}-\hat{\beta}_{j}||\hat{\theta}_{j}X_{i,j}|\leq M|\hat{\theta}_{j}||X_{i,j}|.$ (22)
@@
Let $S^{\prime}:=S\cup\{0\}$. Define $X_{0}\equiv 1$ and $X_{i,0}\equiv 1$ for $1\leq i\leq n$. Combining equations (20), (21), (22) and (Proof of Theorem 1.), we get
@@
Proof. of Proposition 1.
$\sum_{i=1}^{n}(y_{i}-(\theta_{0}+\sum_{j=1}^{p}\hat{\beta}_{0j})-\sum_{j=1}^{p}x_{ij}\hat{\beta}_{j}\theta_{j})^{2}+\sum_{j=1}^{p}|\theta_{j}|$
The result is immediate, since
$|\theta_{j}|=\frac{|\gamma_{j}|}{|\hat{\beta}_{j}|}$
and
[…]
Proof of Theorem 3
Proof.

### v9

@@ 17 Discussion
In this paper we have introduced a novel method for sparse regression that “stacks” univariate regressions using a non-negative version of the lasso. The procedure has interesting properties, namely that the final features weights have the same signs as the univariate coefficients and tend to be larger when the univariate coefficients are large. The test set MSE of the method is often similar to that of the lasso, with substantially smaller support and lower false positive rate.
There are many possible extensions for this work, for example, to random forests, gradient boosting and neural networks. These are topics for future study.
Current versions of the uniLasso package can be found at
https://github.com/trevorhastie/uniLasso and CRAN (R)
https://github.com/sophial05/uni-lasso (Python, written by Sophia Lu).
@@ Acknowledgments
The authors thank two referees and the Editor for helpful comments that improved this manuscript. We also thank Jerome Friedman, Alden Green, Sam Gross, Chris Habron, Lucas Janson, Mert Pilanci, Jonathan Taylor, Ryan Tibshirani and Hui Zou for helpful discussions. We especially thank Ryan for posing and proving theorem 3, and Chris Habron for his contributions of Section 3. S.C. was supported by NSF grants DMS-2113242 and DMS-2153654. T.H. was supported by NIH grant R01 GM134483; R.T. was supported by NIH grant R01 GM134483 and NSF grant 19DMS1208164.
Appendix A Proof of Proposition 1
$\sum_{i=1}^{n}(y_{i}-(\theta_{0}+\sum_{j=1}^{p}\hat{\beta}_{0j})-\sum_{j=1}^{p}x_{ij}\hat{\beta}_{j}\theta_{j})^{2}+\lambda\sum_{j=1}^{p}|\theta_{j}|$
The result is immediate, since
$|\theta_{j}|=\frac{|\gamma_{j}|}{|\hat{\beta}_{j}|}$
and
[…]
Appendix B Proof of theorems in Section 7
We first prove Theorem 1. We will need the following result (see, e.g., (Talagrand, 2010, Theorem A.7.1)).
@@ Appendix B Proof of theorems in Section 7
Choosing $\beta$ sufficiently small makes the upper bound $\leq 2$, allowing us to apply Bernstein’s inequality. A similar argument works for $B_{j}$. For $D_{j}$, take $Z_{i}:=Y_{i}X_{i,j}-\rm E(D_{j})$. Then note that $\rm E(Z_{i})=0$, and for any $t\geq 2|\rm E(D_{j})|$,
@@
There are positive constants $C_{5}$, $C_{6}$ and $C_{7}$ depending only on $C_{0}$, $C_{1}$ and $C_{2}$ such that for any $t\in[0,C_{5}]$ and any $i$ and $j$, $\mathbb{P}(|\hat{\beta}^{-i}_{j}-\beta_{j}|\geq t)$ and $\mathbb{P}(|\hat{\alpha}^{-i}_{j}-\alpha_{j}|\geq t)$ are bounded above by $C_{6}e^{-C_{7}nt^{2}}$.
@@
$\mathbb{P}(|\hat{\beta}^{-i}_{j}-\beta_{j}|\geq t)$ $=\mathbb{P}\biggl{(}\biggl{|}\frac{Q_{1}-Q_{2}Q_{3}-(Q_{4}-Q_{3}^{2})\beta_{j}}{Q_{4}-Q_{3}^{2}}\biggr{|}\geq t\biggr{)}$
@@
Choosing $s$ to solve $K_{5}s+K_{6}s^{2}=C_{0}t/2$, and combining equations (11), (13) and (14) yields the proof of the claimed concentration inequality for $\hat{\beta}^{-i}_{j}$. The proof for $\hat{\alpha}^{-i}_{j}$ is similar. We omit the details. ∎
@@
$\frac{1}{n}\sum_{i=1}^{n}(\hat{Y}_{i,j}-\hat{\alpha}_{j})^{2}$ $\leq\frac{2}{n}\sum_{i=1}^{n}(\hat{Y}_{i,j}-\hat{\alpha}^{-i}_{j})^{2}+\frac{2}{n}\sum_{i=1}^{n}(\hat{\alpha}^{-i}_{j}-\hat{\alpha}_{j})^{2}$
$\leq\frac{2}{n}\sum_{i=1}^{n}\hat{\beta}^{2}_{i,j}X_{i,j}^{2}+\frac{4}{n}\sum_{i=1}^{n}(\hat{\alpha}^{-i}_{j}-\alpha_{j})^{2}+4(\hat{\alpha}_{j}-\alpha_{j})^{2}$
$\leq\biggl{(}\frac{2}{n}\sum_{i=1}^{n}X_{i,j}^{2}\biggr{)}\max_{1\leq i\leq n}(\hat{\beta}^{-i}_{j})^{2}+4\max\{(\hat{\alpha}_{j}-\alpha_{j})^{2},\max_{1\leq i\leq n}(\hat{\alpha}^{-i}_{j}-\alpha_{j})^{2}\}.$
@@
because otherwise, we can slightly perturb $\hat{\theta}_{j}$ while maintaining the non-negativity constraint and decreasing the value of $L$. By inequality (16), this implies that if $\hat{\theta}_{j}>0$, then we must have
@@
$M:=\max_{1\leq i\leq n,\,j\in S}(|\hat{\beta}^{-i}_{j}-\beta_{j}|+|\hat{\beta}_{j}-\beta_{j}|).$
@@
$|\hat{\theta}_{j}(\hat{Y}_{i,j}-\hat{\alpha}_{j})-\hat{\gamma}_{j}X_{i,j}|$ $=|\hat{\theta}_{j}\hat{\beta}^{-i}_{j}X_{i,j}-\hat{\theta}_{j}\hat{\beta}_{j}X_{i,j}|$
$=|\hat{\beta}^{-i}_{j}-\hat{\beta}_{j}||\hat{\theta}_{j}X_{i,j}|\leq M|\hat{\theta}_{j}||X_{i,j}|.$ (21)
@@
Let $S^{\prime}:=S\cup\{0\}$. Define $X_{0}\equiv 1$ and $X_{i,0}\equiv 1$ for $1\leq i\leq n$. Combining equations (19), (20), (21) and (B), we get
@@
Proof of Theorem 3.